OpenStudy (anonymous):

Factor completely, then place the factors in the proper location on the grid. 2a2 + 2b2 - 5ab

OpenStudy (anonymous):

Assuming "b2" is referring to functions | Use as a unit instead Assuming "ab" is concatenated variables | Use as a unit instead Input: More formsStep-by-step solution (2 a-b) (a-2 b) a (2 a-5 b)+2 b^2 2 a^2+b (2 b-5 a) Real roots: b = a/2 b = 2 a

OpenStudy (anonymous):

thanks